# Implicit Language Models are RNNs: Balancing Parallelization and Expressivity

### v2

@@ Abstract
State-space models (SSMs) and transformers dominate the language modeling landscape. However, they are constrained to a lower computational complexity than classical recurrent neural networks (RNNs), limiting their expressivity. In contrast, RNNs lack parallelization during training, raising fundamental questions about the trade off between parallelization and expressivity. We propose implicit SSMs, which iterate a transformation until convergence to a fixed point. Theoretically, we show that implicit SSMs implement the non-linear state-transitions of RNNs. Empirically, we find that only approximate fixed-point convergence suffices, enabling the design of a scalable training curriculum that largely retains parallelization, with full convergence required only for a small subset of tokens. Our approach demonstrates superior state-tracking capabilities on regular languages, surpassing transformers and SSMs. We further scale implicit SSMs to natural language reasoning tasks and pretraining of large-scale language models up to 1.3B parameters on 207B tokens–representing, to our knowledge, the largest implicit model trained to date. Notably, our implicit models outperform their explicit counterparts on standard benchmarks.
@@ 1 Introduction
Figure 1: Left: Minimum layers required to solve the $S_{5}$ word problem, a theoretically hard formalization of state tracking, for different sequence lengths. Right: Length generalization for Mamba2 and our implicit Mamba2 trained on $L=32$ and extrapolated up to $L=128$. Bottom: Scaling of language models pretrained on $207$B tokens of the deduplicated Pile.
@@ 2.2 Limitations of Transformers and SSMs
Efficient parallelization is one of the central features enabling transformers and SSMs to scale to large machine learning problems such as language modeling. Parallel circuits, however, face fundamental trade-offs regarding the class of problems that they can address. In particular, transformers and SSMs theoretically fail to recognize certain regular languages, or equivalently, to simulate FSMs (Merrill et al., 2022, 2024). Empirical studies have confirmed that neither of the models are capable of learning the algorithms constituting certain regular languages (Bhattamishra et al., 2020; Sarrof et al., 2024). By contrast, the sequential nature of RNNs allows them to express all regular languages (Merrill, 2019). A detailed discussion is given in Section A.1.
@@ 2.3 Deep Equilibrium Models
A range of methods have been proposed to efficiently compute $\frac{\partial\mathcal{L}}{\partial\theta}=\frac{\partial\Phi}{\partial\theta}\frac{\partial\mathcal{L}}{\partial z^{\ast}}$ using Equation (4) (Bai et al., 2019; Geng et al., 2021). Here, we employ the Phantom Gradient approach of Geng et al. (2021) (see in the Appendix Figure 5). The method is based on solving a smoothed version of the fixed point equation (3) combined with a finite truncation of the von Neumann series of the Jacobian-vector-product in (4) given as
@@ 3.1 Implicit State-space Models
The linear recurrence of SSMs shown in equation (1) cannot resolve elaborate sequential problems (Merrill et al., 2024). Here, we propose to exploit self-iterations along the depth of neural networks to close the expressivity gap between SSMs and RNNs. Following the DEQ paradigm (see Section 2.3), we implicitly define a model via the fixed points of a SSM. Introducing the iteration variable $z_{t}^{(s)}\in\mathbb{R}^{d_{\text{out}}}$ to equations (1) and (2) yields the fixed point iteration
@@
$z_{t}^{\left(s\right)}$ $=f_{\theta}\left(z_{t}^{\left(s-1\right)},h_{t}^{\left(s\right)},x_{t}\right)\,,$ (7)
where ${z_{t}^{(0)}=0}$ for ${t=0,\dots,T}$ and ${h_{0}^{(s)}=0}$ for ${s=0,\dots,S}$ respectively. The output $z^{(s)}_{t}$ of $f_{\theta}$ is fed back to the self-iteration until (approximate) convergence to a fixed-point. Note, how this adds a new dependency to the functions $\Lambda$ and $u$ in equation (6) that is not present in equation (1). Notably, this minor technical change leads to fundamental differences between explicit SSMs and the implicit SSM defined above.
Computing the output, as well as the gradient, of our implicit SSM requires to iterate the two loops defining equations (6) and (7): A loop $t=1,\dots,T$ over the sequence dimension, and a loop $s=0,\dots,S$ to find the fixed point. The two loops give rise to two modes of evaluation visualized in Figure 2. The simultaneous mode simultaneously finds the fixed points for all $t$ (see Figure 2A), and exploits parallelization strategies for SSMs (Dao & Gu, 2024). The sequential mode resolves the $s$ and $t$ loops in the transpose order, and processes sequences sequentially just like classical SSMs or RNNs (see Figure 2B). While the simultaneous mode allows for highly parallel training, the sequential mode enables efficient inference at constant memory, e.g. for language generation. For both modes, equation (6) in the limit ${s\rightarrow\infty}$ reads
where ${z_{t}^{*}=\lim_{s\rightarrow\infty}z_{t}^{(s)}}$ and ${h_{t}^{*}=\lim_{s\rightarrow\infty}h_{t}^{(s)}}$ denote the fixed points. The fixed point $z_{t}^{*}$ depends on $h_{t}^{*}$, and hence by equation (7) on $h_{t-1}^{*}$. Notably, our self-iteration introduces a non-linear dependency to the originally linear recurrence (1) via the functions $\Lambda$ and $u$. Thereby, our implicit SSM inherits one of the crucial properties of RNNs, as formalized next.
@@
Consider an implicit SSM defined by equations (6) and (7). Then the transition function $h_{t-1}^{*}\mapsto h_{t}^{*}$ defined by equation (8) is non-linear and non-diagonal, i.e. each hidden state $h_{t}^{*}$ is a non-linear function of the previous hidden state $h_{t-1}^{*}$. Consequently, the state-to-state Jacobian is a non-diagonal operator.
Proof: We refer the reader to Appendix B
As discussed in Section 2.2, non-linear RNNs surpass transformers and linear SSMs in terms of circuit complexity. By the above construction, our implicit SSM appears to exhibit the favourable computational properties of RNNs, lifting the illusion of state in linear SSMs  (Merrill et al., 2024). Furthermore, the gradients of a fixed point iteration depend solely on the fixed point, and not on the path to the fixed point, by the implicit function theorem. This suggests that both modes resolving the two for loops yield functionally equivalent fixed points.
@@ 3.2 Implicit Transformers
Similar to implicit SSMs, one can define an implicit transformer model (Bai et al., 2019) as
$z^{(s)}_{t}=\operatorname{LN}(\operatorname{FFN}(\operatorname{LN}(\operatorname{Attn}(z^{(s-1)}_{t}W_{QKV}+x_{t}W_{\text{inp}})))),$
where $W_{QKV}\in\mathbb{R}^{d\times 3d}$ produces the $Q$, $K$, $V$ for the multi-head self-attention (Attn), FFN denotes a feed-forward block, LN stands for layer normalization, and $W_{\text{inp}}\in\mathbb{R}^{d\times 3d}$ is the input projection. Conventional transformers, with their finite number of layers, cannot learn certain formal languages outside of the TC0 circuit complexity class (Merrill et al., 2022; Strobl et al., 2024). However, chain of thought (CoT) models (Wei et al., 2022) bypass this restriction by using an adaptive compute budget through recursive generation of intermediate tokens (Merrill & Sabharwal, ). Implicit transformers (Bai et al., 2019) utilize an adaptive compute budget differently, using fixed-point iterations that can be interpreted as sequences of latent thoughts (Hao et al., 2024), undergoing non-linear updates similar to a non-linear RNN’s hidden state.
@@ 4 Implicit SSMs Adapt to Hard Languages
The Illusion of State (Merrill et al., 2024) reveals that SSMs cannot simulate arbitrary finite state machines. A hard state tracking problem in the sense that all state tracking problems can be reduced to it is given by the word problem for the symmetric group $S_{5}$ (Barrington, 1989). The word problem for a monoid $\left(M,\circ\right)$ is to resolve arbitrary length products of the form ${\hat{m}=m_{1}\cdot m_{2}\circ\dots\circ m_{k}}$ for ${m_{1},m_{2},\dots,m_{k}\in M,k\in\mathbb{N}}$. A comprehensive introduction to the word problem and our particular learning setting is provided in Section D.1.
@@
We will identify minimally sequential models that parallelize to a high degree and still capture all non-solvable transitions in a language. Therefore, we apply our construction of a word problem above to mix tokens from simple languages with tokens from non-solvable hard languages. This section studies a word problem over $M=M^{a}\times A5$, where $M^{a}$ is a simple aperiodic monoid with four elements and $A_{5}\subset S_{5}$ is the alternating group over 5 elements, the smallest non-solvable subgroup of $S_{5}$. For details on the learning problem, we refer the reader to Section D.1.
@@ 5 Implicit Large Language Models
Model Average
$\downarrow$
Mamba2 25.37
Mamba2 15.06
Mamba2(4+1)-ours 9.53
Mamba2(24+4)-ours 9.63
Figure 4: Length extrapolation performance on the the test split of the D-Pile of the original 1.3B Mamba2, our Mamba2, and our implicit Mamba2 with (4+1) and (24+4) self-iterations. Shaded gray area shows the in-distribution length. Left: Per token perplexities at different lengths. Right: The average perplexity of tokens for a context length of 16 3841638416\,38416 384.
All implicit models in our study were trained on sequences of 2048 tokens. To assess their capability for length extrapolation, we evaluated the implicit models on the test split of the D-Pile, which was packed with longer sequences consisting of 4096, 8192, and 16384 tokens. We compared these results with the baseline Mamba2 and Mamba2 in Figure 4, where the per-token perplexities are reported. For the average perplexity at 16384 and other lengths, refer to the table in Figure 4 and Table 3 in the Appendix. The implicit Mamba2 models maintain their perplexity as sequence length increases, whereas the baseline Mamba2 models exhibit an increase in perplexity with longer sequences.
@@
Table 1: Comparison of test set perplexity and downstream performance. We compare our implicit models, which have 4 self-iteration steps and 1 phantom gradient step (denoted as 4+1), and those with 24/32 self-iteration steps and 4 phantom gradient steps (denoted as 24+4/32+4), with our baseline models Mamba2 and Llama. These baseline models as well as the implicit models are trained on 207B tokens from the D-Pile dataset and range in size from 130M to 1.3B parameters. For further comparison, we include the original Mamba2  (Dao & Gu, 2024) (trained on 300B tokens of the Pile) and the Llama (trained on 300B tokens of the SlimPajama) from (Beck et al., 2024). The best performing model for each type is highlighted in bold, and the second-best is underlined.
@@
Model Dataset/Tokens (B) D-Pile LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande OpenbookQA Average
$\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
130M Mamba2 Pile/300 13.72 16.83 0.4388 0.3525 0.6496 0.4739 0.2423 0.5233 0.306 0.4266
Mamba2 D-Pile/207 13.05 18.51 0.4116 0.3527 0.6572 0.4815 0.2372 0.5130 0.300 0.4219
Mamba2(4+1)-ours D-Pile/207 13.76 18.58 0.4118 0.3628 0.6485 0.4537 0.2287 0.5107 0.288 0.4149
Mamba2(24+4)-ours D-Pile/207 12.86 18.03 0.4174 0.3673 0.6496 0.4604 0.2372 0.5178 0.290 0.4200
Llama SlimPajama/300 - 39.21 0.3154 0.3409 0.6545 0.4533 0.2363 0.5067 - 0.4178
Llama D-Pile/207 12.77 17.08 0.4297 0.3513 0.6540 0.4794 0.2440 0.5122 0.280 0.4215
Llama (4+1)-ours D-Pile/207 12.73 15.54 0.4518 0.3706 0.6447 0.4823 0.2372 0.5391 0.290 0.4308
Llama (32+4)-ours D-Pile/207 11.73 13.39 0.4801 0.3958 0.6676 0.4886 0.2355 0.5304 0.298 0.4423
350M Mamba2 Pile/300 10.55 8.00 0.5593 0.4692 0.7046 0.5476 0.2671 0.5564 0.324 0.4897
Mamba2 D-Pile/207 10.18 8.96 0.5333 0.4653 0.6942 0.5526 0.2696 0.5320 0.306 0.4790
Mamba2(4+1)-ours D-Pile/207 10.02 8.79 0.5457 0.4684 0.6899 0.5358 0.2696 0.5162 0.308 0.4762
Mamba2(24+4)-ours D-Pile/207 9.70 8.26 0.5575 0.4792 0.7040 0.5484 0.2688 0.5351 0.316 0.487
Llama SlimPajama/300 - 15.73 0.4419 0.4445 0.6915 0.5223 0.2628 0.5359 - 0.4832
Llama D-Pile/207 10.30 8.37 0.5624 0.4537 0.6844 0.5476 0.2577 0.5541 0.318 0.4826
Llama (4+1)-ours D-Pile/207 9.66 7.03 0.5898 0.5030 0.7024 0.5539 0.2611 0.5572 0.314 0.4973
Llama (32+4)-ours D-Pile/207 9.43 7.04 0.5956 0.5114 0.7078 0.5244 0.2705 0.5722 0.320 0.5003
760M Mamba2 Pile/300 9.23 5.86 0.6167 0.5492 0.7198 0.6103 0.2850 0.6030 0.362 0.5351
Mamba2 D-Pile/207 8.98 6.24 0.6125 0.5418 0.7231 0.6044 0.2858 0.5777 0.338 0.5262
Mamba2(4+1)-ours D-Pile/207 8.60 6.15 0.6117 0.5569 0.7296 0.6077 0.3140 0.5509 0.336 0.5295
Mamba2(24+4)-ours D-Pile/207 8.35 5.90 0.6191 0.5698 0.7334 0.6090 0.3131 0.5730 0.338 0.5365
Llama SlimPajama/300 - 9.90 0.5141 0.5216 0.7095 0.5648 0.2875 0.5667 - 0.5274
Llama D-Pile/207 8.88 5.77 0.6375 0.5448 0.7171 0.5905 0.2816 0.6054 0.338 0.5307
Llama (4+1)-ours D-Pile/207 8.27 5.15 0.6524 0.5853 0.7312 0.6052 0.3097 0.5967 0.356 0.5481
Llama (32+4)-ours D-Pile/207 7.90 4.82 0.6703 0.5995 0.7416 0.6187 0.3012 0.5991 0.344 0.5535
1.3B Mamba2 Pile/300 8.40 5.02 0.6559 0.5995 0.7378 0.6418 0.3319 0.6117 0.378 0.5652
Mamba2 D-Pile/207 8.28 5.12 0.6456 0.5939 0.7416 0.6145 0.3123 0.6117 0.352 0.5531
Mamba2(4+1)-ours D-Pile/207 7.97 5.21 0.6383 0.6136 0.7437 0.6343 0.3302 0.5746 0.354 0.5555
Mamba2(24+4)-ours D-Pile/207 7.70 4.99 0.6489 0.6267 0.7416 0.6423 0.3336 0.5888 0.352 0.5620
Llama SlimPajama/300 - 7.23 0.5744 0.5781 0.7312 0.6279 0.3174 0.5904 - 0.5699
Llama D-Pile/207 7.99 4.95 0.6569 0.5936 0.7432 0.6385 0.3217 0.6062 0.352 0.5589
Llama (4+1)-ours D-Pile/207 7.66 4.40 0.6852 0.6397 0.7448 0.6338 0.3396 0.6575 0.360 0.5801
Llama (32+4)-ours D-Pile/207 7.24 4.24 0.6901 0.6583 0.7465 0.6654 0.3601 0.6401 0.364 0.5892

### v3

@@ Abstract
State-space models (SSMs) and transformers dominate the language modeling landscape. However, they are constrained to a lower computational complexity than classical recurrent neural networks (RNNs), limiting their expressivity. In contrast, RNNs lack parallelization during training, raising fundamental questions about the trade off between parallelization and expressivity. We propose implicit SSMs, which iterate a transformation until convergence to a fixed point. Theoretically, we show that implicit SSMs implement the non-linear state-transitions of RNNs. Empirically, we find that only approximate fixed-point convergence suffices, enabling the design of a scalable training curriculum that largely retains parallelization, with full convergence required only for a small subset of tokens. Our approach demonstrates superior state-tracking capabilities on regular languages, surpassing transformers and SSMs. We further scale implicit SSMs to natural language reasoning tasks and pretraining of large-scale language models up to 1.3B parameters on 207B tokens–representing, to our knowledge, the largest implicit model trained to date. Notably, our implicit models outperform their explicit counterparts on standard benchmarks. Our code is publicly available at github.com/microsoft/implicit_languagemodels.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Top Left: Minimum layers required to solve the $S_{5}$ word problem, a theoretically hard formalization of state tracking, for different sequence lengths. Top Right: Length generalization for Mamba2 and our implicit Mamba2 trained on $L=32$ and extrapolated up to $L=128$. Center: Downstream task accuracy for a range of self-iterations at test-time for our implicit $1.3$B parameter Mamba2 and Llama3 models. Dashed lines indicate explicit models’ accuracy. Bottom: Scaling of language models pretrained on $207$B tokens of the deduplicated Pile.
@@ 2.2 Limitations of Transformers and SSMs
Efficient parallelization is one of the central features enabling transformers and SSMs to scale to large machine learning problems such as language modeling. Parallel circuits, however, face fundamental trade-offs regarding the class of problems that they can address. In particular, transformers and SSMs theoretically fail to recognize certain regular languages, or equivalently, to simulate FSMs (Merrill et al., 2022, 2024). Empirical studies have confirmed that neither of the models are capable of learning the algorithms constituting certain regular languages (Bhattamishra et al., 2020; Sarrof et al., 2024). By contrast, the sequential nature of RNNs allows them to express all regular languages (Merrill, 2019). A detailed discussion is given in Section B.1.
@@ 2.3 Deep Equilibrium Models
A range of methods have been proposed to efficiently compute $\frac{\partial\mathcal{L}}{\partial\theta}=\frac{\partial\Phi}{\partial\theta}\frac{\partial\mathcal{L}}{\partial z^{\ast}}$ using Equation 4 (Bai et al., 2019; Geng et al., 2021). Here, we employ the Phantom Gradient approach of Geng et al. (2021) (see in the Appendix Figure 5). The method is based on solving a smoothed version of the fixed point Equation 3 combined with a finite truncation of the von Neumann series of the Jacobian-vector-product in (4) given as
@@ 3.1 Implicit State-space Models
The linear recurrence of SSMs shown in Equation 1 cannot resolve elaborate sequential problems (Merrill et al., 2024). Here, we propose to exploit self-iterations along the depth of neural networks to close the expressivity gap between SSMs and RNNs. Following the DEQ paradigm (see Section 2.3), we implicitly define a model via the fixed points of a SSM. Introducing the iteration variable $z_{t}^{(s)}\in\mathbb{R}^{d_{\text{out}}}$ to Equations 1 and 2 yields the fixed point iteration
@@
$z_{t}^{\left(s\right)}$ $=f_{\theta}\left(z_{t}^{\left(s-1\right)},h_{t-1}^{\left(s\right)},x_{t}\right)\,,$ (7)
where ${z_{t}^{(0)}=0}$ for ${t=0,\dots,T}$ and ${h_{0}^{(s)}=0}$ for ${s=0,\dots,S}$ respectively. In practice, we rely on the Mamba2 (Dao & Gu, 2024) architecture and inject the expanded input $x_{t}$ additively after the initial projection.
Computing the fixed-points and the gradient around the fixed points of Equations 6 and 7 requires to iterate the two loops $t=1,\dots,T$ along the input sequence and $s=0,\dots,S$ to converge to fixed points along depth. Figure 2 visualizes how these two loops give rise to two distinct modes of evaluation. The simultaneous mode simultaneously finds the fixed points for all $t$ (see Figure 2A), and exploits parallelization strategies for SSMs (Dao & Gu, 2024). The sequential mode resolves the $s$ and $t$ loops in the transpose order, and processes sequences sequentially just like classical SSMs or RNNs (see Figure 2B). While the simultaneous mode allows for highly parallel training, the sequential mode enables efficient inference at constant memory, e.g. for language generation. The emergence of these two strategies for finding the fixed points sparks the question if they are interchangeable as visualized in Figure 6. Can a model be trained in the simultaneous mode and deployed for inference in the sequential mode? This question is answered in Section 5 and Figure 2 for real language models.
For both modes, Equations 6 and 7 in the limit ${s\rightarrow\infty}$ read
$h_{t}^{*}$ $=\Lambda\left(z_{t}^{*},x_{t}\right)h_{t-1}^{*}+u\left(z_{t}^{*},x_{t}\right)$ (8)
$z_{t}^{*}$ $=f_{\theta}\left(z_{t}^{*},h_{t-1}^{*},x_{t}\right)\,,$ (9)
where ${z_{t}^{*}=\lim_{s\rightarrow\infty}z_{t}^{(s)}}$ and ${h_{t}^{*}=\lim_{s\rightarrow\infty}h_{t}^{(s)}}$ denote the fixed points. Equations 8 and 9 couple the outputs $z_{t}^{*}$ with the states $h_{t}^{*}$ via the non-linear functions $f_{\theta},\Lambda$ and $u$. Notably, what appears as a minor technical modification of the original state-space model in Equations 1 and 2 leads to a fundamentally enhanced expressivity of the recurrent model.
@@
An implicit SSM defined by Equations 8 and 9 with generic weights yields a non-linear and non-diagonal state-to-state transition function $h_{t-1}^{*}\mapsto h_{t}^{*}$. If further ${z_{t}^{*}-f_{\theta}\left(z_{t}^{*},h_{t-1}^{*},x_{t}\right)=0}$ is non-singular w.r.t. $z_{t}^{*}$, then there exists a continuously differentiable function $\varphi$ such that $z_{t}^{*}=\varphi\left(h_{t}^{*},x_{t},\theta\right)$. In this case, the state-to-state Jacobian is given by
$\left.\frac{\mathrm{d}h_{t}^{*}}{\mathrm{d}h_{t-1}^{*}}\right|_{z_{t}^{*},x_{t},\theta}=\Lambda\left(z_{t}^{*},x_{t}\right)$ $+\frac{\partial\Lambda}{\partial z_{t}^{*}}\frac{\partial\varphi}{\partial h_{t-1}^{*}}\mathrm{diag}\left(h_{t-1}^{*}\right)$
Proof: We refer the reader to Section A.2 for the proof and an extended statement of the theorem. A numerical check of Equation 10 is provided in Figure 7. ∎
As discussed in Section 2.2, non-linear RNNs surpass transformers and linear SSMs in terms of circuit complexity. By the above construction, our implicit SSM exhibits the favourable computational properties of RNNs, lifting the illusion of state in linear SSMs  (Merrill et al., 2024). Furthermore, the gradients of a fixed point iteration depend solely on the fixed point, and not on the path to the fixed point, by the implicit function theorem. This suggests that both modes resolving the two for loops yield functionally equivalent fixed points.
@@ 3.2 Implicit Transformers
Similar to implicit SSMs, one can define an implicit transformer model (Bai et al., 2019) by injecting the input $x_{t}$ into the attention operator as $\operatorname{Attn}(z_{t}W_{QKV}+x_{t}W_{\text{inp}})$, where $W_{QKV}\in\mathbb{R}^{d\times 3d}$ produces the $Q$, $K$, $V$ for the multi-head self-attention (Attn), and $W_{\text{inp}}\in\mathbb{R}^{d\times 3d}$ is the input projection. Conventional transformers, with their finite number of layers, cannot learn certain formal languages outside of the TC0 circuit complexity class (Merrill et al., 2022; Strobl et al., 2024). However, chain of thought (CoT) models (Wei et al., 2022) bypass this restriction by using an adaptive compute budget through recursive generation of intermediate tokens (Merrill & Sabharwal, 2024). Implicit transformers (Bai et al., 2019) utilize an adaptive compute budget differently, using fixed-point iterations that can be interpreted as sequences of latent thoughts (Hao et al., 2024), undergoing non-linear updates similar to a non-linear RNN’s hidden state.
@@ 4 Implicit SSMs Adapt to Hard Languages
The illusion of state (Merrill et al., 2024) reveals that constant depth SSMs cannot simulate arbitrary finite state machines. A hard state tracking problem in the sense that all state tracking problems can be reduced to it is given by the word problem for the symmetric group $S_{5}$ (Barrington, 1989). The word problem for a monoid $\left(M,\circ\right)$ is to resolve arbitrary length products of the form ${\hat{m}=m_{1}\cdot m_{2}\circ\dots\circ m_{k}}$ for ${m_{1},m_{2},\dots,m_{k}\in M,k\in\mathbb{N}}$. A comprehensive introduction to the word problem and our particular learning setting is provided in Appendix C.
@@
We will identify minimally sequential models that parallelize to a high degree and still capture all non-solvable transitions in a language. Therefore, we apply our construction of a word problem above to mix tokens from simple languages with tokens from non-solvable hard languages. This section studies a word problem over $M=M^{a}\times A5$, where $M^{a}$ is a simple aperiodic monoid with four elements and $A_{5}\subset S_{5}$ is the alternating group over 5 elements, the smallest non-solvable subgroup of $S_{5}$. For details on the learning problem, we refer the reader to Appendix C.
@@ 5 Implicit Large Language Models
Figure 4: Length extrapolation performance of the per token perplexity on the the test split of the D-Pile of the original 1.3B Mamba2, our Mamba2, and our implicit Mamba2 with (4+1) and (24+4) self-iterations. Shaded gray area shows the in-distribution length.
All implicit models in our study were trained on sequences of 2048 tokens. To assess their capability for length extrapolation, we evaluated the implicit models on the test split of the D-Pile, which was packed with longer sequences consisting of 4096, 8192, and 16384 tokens. We compared these results with the baseline Mamba2 and Mamba2 in Figure 4, where the per-token perplexities are reported. Numerical values for all models at a range of token positions are reported in Table 7 in the Appendix. In comparison with the baseline Mamba2 models, our implicit Mamba2 models demonstrate stronger robustness on longer sequences.
@@
Table 1: Comparison of test set perplexity and downstream performance. We compare our implicit models, which have 4 self-iteration steps and 1 phantom gradient step (denoted as 4+1), and those with 24/32 self-iteration steps and 4 phantom gradient steps (denoted as 24+4/32+4), with our baseline models Mamba2 and Llama. These baseline models as well as the implicit models are trained on 207B tokens from the D-Pile dataset and range in size from 130M to 1.3B parameters. For further comparison, we include the original Mamba2  (Dao & Gu, 2024) (trained on 300B tokens of the Pile) and the Llama (trained on 300B tokens of the SlimPajama) from (Beck et al., 2024). The best performing model for each type is highlighted in bold, and the second-best is underlined. The training wall-clock time (WCT) is provided for implicit models relative to their explicit counterpart. For $(24+4)$- and $(32+4)$-iterated models, we provided the weighted average measured over the entire curriculum. We omit WCTs for the 370M-model series since the runs were carried out on various node sizes, making direct comparison difficult.
@@
Model Dataset/ Rel. WCT D-Pile LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande OpenbookQA Average
Tokens (B) $\nicefrac{{\textrm{Impl}}}{{\textrm{Expl}}}$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
130M Mamba2 Pile/300 - 13.72 16.83 0.4388 0.3525 0.6496 0.4739 0.2423 0.5233 0.306 0.4266
Mamba2 D-Pile/207 1 13.05 18.51 0.4116 0.3527 0.6572 0.4815 0.2372 0.5130 0.300 0.4219
Mamba2(4+1)-ours D-Pile/207 1.97 13.76 18.58 0.4118 0.3628 0.6485 0.4537 0.2287 0.5107 0.288 0.4149
Mamba2(24+4)-ours D-Pile/207 3.27 12.86 18.03 0.4174 0.3673 0.6496 0.4604 0.2372 0.5178 0.290 0.4200
Llama SlimPajama/300 - - 39.21 0.3154 0.3409 0.6545 0.4533 0.2363 0.5067 - 0.4178
Llama D-Pile/207 1 12.77 17.08 0.4297 0.3513 0.6540 0.4794 0.2440 0.5122 0.280 0.4215
Llama (4+1)-ours D-Pile/207 0.90 12.73 15.54 0.4518 0.3706 0.6447 0.4823 0.2372 0.5391 0.290 0.4308
Llama (32+4)-ours D-Pile/207 2.08 11.73 13.39 0.4801 0.3958 0.6676 0.4886 0.2355 0.5304 0.298 0.4423
370M Mamba2 Pile/300 - 10.55 8.00 0.5593 0.4692 0.7046 0.5476 0.2671 0.5564 0.324 0.4897
Mamba2 D-Pile/207 - 10.18 8.96 0.5333 0.4653 0.6942 0.5526 0.2696 0.5320 0.306 0.4790
Mamba2(4+1)-ours D-Pile/207 - 10.02 8.79 0.5457 0.4684 0.6899 0.5358 0.2696 0.5162 0.308 0.4762
Mamba2(24+4)-ours D-Pile/207 - 9.70 8.26 0.5575 0.4792 0.7040 0.5484 0.2688 0.5351 0.316 0.487
Llama SlimPajama/300 - - 15.73 0.4419 0.4445 0.6915 0.5223 0.2628 0.5359 - 0.4832
Llama D-Pile/207 - 10.30 8.37 0.5624 0.4537 0.6844 0.5476 0.2577 0.5541 0.318 0.4826
Llama (4+1)-ours D-Pile/207 - 9.66 7.03 0.5898 0.5030 0.7024 0.5539 0.2611 0.5572 0.314 0.4973
Llama (32+4)-ours D-Pile/207 - 9.43 7.04 0.5956 0.5114 0.7078 0.5244 0.2705 0.5722 0.320 0.5003
760M Mamba2 Pile/300 - 9.23 5.86 0.6167 0.5492 0.7198 0.6103 0.2850 0.6030 0.362 0.5351
Mamba2 D-Pile/207 1 8.98 6.24 0.6125 0.5418 0.7231 0.6044 0.2858 0.5777 0.338 0.5262
Mamba2(4+1)-ours D-Pile/207 2.05 8.60 6.15 0.6117 0.5569 0.7296 0.6077 0.3140 0.5509 0.336 0.5295
Mamba2(24+4)-ours D-Pile/207 3.43 8.35 5.90 0.6191 0.5698 0.7334 0.6090 0.3131 0.5730 0.338 0.5365
Llama SlimPajama/300 - - 9.90 0.5141 0.5216 0.7095 0.5648 0.2875 0.5667 - 0.5274
Llama D-Pile/207 1 8.88 5.77 0.6375 0.5448 0.7171 0.5905 0.2816 0.6054 0.338 0.5307
Llama (4+1)-ours D-Pile/207 1.64 8.27 5.15 0.6524 0.5853 0.7312 0.6052 0.3097 0.5967 0.356 0.5481
Llama (32+4)-ours D-Pile/207 2.97 7.90 4.82 0.6703 0.5995 0.7416 0.6187 0.3012 0.5991 0.344 0.5535
1.3B Mamba2 Pile/300 - 8.40 5.02 0.6559 0.5995 0.7378 0.6418 0.3319 0.6117 0.378 0.5652
Mamba2 D-Pile/207 1 8.28 5.12 0.6456 0.5939 0.7416 0.6145 0.3123 0.6117 0.352 0.5531
Mamba2(4+1)-ours D-Pile/207 1.91 7.97 5.21 0.6383 0.6136 0.7437 0.6343 0.3302 0.5746 0.354 0.5555
Mamba2(24+4)-ours D-Pile/207 3.19 7.70 4.99 0.6489 0.6267 0.7416 0.6423 0.3336 0.5888 0.352 0.5620
Llama SlimPajama/300 - - 7.23 0.5744 0.5781 0.7312 0.6279 0.3174 0.5904 - 0.5699
Llama D-Pile/207 1 7.99 4.95 0.6569 0.5936 0.7432 0.6385 0.3217 0.6062 0.352 0.5589
Llama (4+1)-ours D-Pile/207 1.96 7.66 4.40 0.6852 0.6397 0.7448 0.6338 0.3396 0.6575 0.360 0.5801
Llama (32+4)-ours D-Pile/207 3.91 7.24 4.24 0.6901 0.6583 0.7465 0.6654 0.3601 0.6401 0.364 0.5892